# Classic GNNs are Strong Baselines: Reassessing GNNs for Node Classification

### v1

@@ Abstract
Graph Transformers (GTs) have recently emerged as popular alternatives to traditional message-passing Graph Neural Networks (GNNs), due to their theoretically superior expressiveness and impressive performance reported on standard node classification benchmarks, often significantly outperforming GNNs. In this paper, we conduct a thorough empirical analysis to reevaluate the performance of three classic GNN models (GCN, GAT, and GraphSAGE) against GTs. Our findings suggest that the previously reported superiority of GTs may have been overstated due to suboptimal hyperparameter configurations in GNNs. Remarkably, with slight hyperparameter tuning, these classic GNN models achieve state-of-the-art performance, matching or even exceeding that of recent GTs across 17 out of the 18 diverse datasets examined. Additionally, we conduct detailed ablation studies to investigate the influence of various GNN configurations—such as normalization, dropout, residual connections, network depth, and jumping knowledge mode—on node classification performance. Our study aims to promote a higher standard of empirical rigor in the field of graph machine learning, encouraging more accurate comparisons and evaluations of model capabilities. Our implementation is available at https://github.com/LUOyk1999/tunedGNN.
@@ 1 Introduction
Node classification is a fundamental task in graph machine learning, with high-impact applications across many fields such as social network analysis, bioinformatics, and recommendation systems. Graph Neural Networks (GNNs)  [20, 28, 65, 72, 49, 33, 8, 76, 18, 9, 51, 4, 73, 56, 10, 52, 53, 66, 77, 37] have emerged as a powerful class of models for tackling the node classification task. GNNs operate by iteratively aggregating information from a node’s neighbors, a process known as message passing [19], leveraging both the graph structure and node features to learn useful node representations for classification. While GNNs have achieved notable success, studies have identified several limitations, including over-smoothing [35], over-squashing [1], lack of sensitivity to heterophily [81], and challenges in capturing long-range dependencies [11].
@@
This prompts us to reconsider: Could the potential of message-passing GNNs for node classification have been previously underestimated? While prior research has addressed this issue to some extent [24, 14, 68, 44, 54], these studies have limitations in terms of scope and comprehensiveness, including a restricted number and diversity of datasets, as well as an incomplete examination of hyperparameters. In this study, we comprehensively reassess the performance of GNNs for node classification, utilizing three classic GNN models—GCN [28], GAT [64], and GraphSAGE [20]—across 18 real-world benchmark datasets that include homophilous, heterophilous, and large-scale graphs. We examine the influence of key hyperparameters on GNN training, including normalization [2, 26], dropout [63], residual connections [21], network depth, and the utilization of the jumping knowledge mode [73]. We summarize the key findings in our empirical study as follows:
@@
  • Our ablation studies have yielded valuable insights into GNN hyperparameters for node classification. We demonstrate that (1) normalization is essential for large-scale graphs; (2) dropout consistently proves beneficial; (3) residual connections can significantly enhance performance, especially on heterophilous graphs; (4) GNNs on heterophilous graphs tend to perform better with deeper layers; and (5) jumping knowledge connections have minimal impact on large-scale graphs.
@@ 3 Key Hyperparameters for Training GNNs
In this section, we present an overview of the key hyperparameters for training GNNs, including normalization, dropout, residual connections, network depth, and “jumping knowledge” mode. The first four hyperparameters are widely utilized across different types of neural networks to improve model performance, while the last one is specific to GNNs.
@@
$\boldsymbol{h}_{v}^{l}=\text{Dropout}(\sigma(\text{Norm}(\boldsymbol{h}_{u}^{l-1}\boldsymbol{W_{r}}^{l}+\sum_{u\in\mathcal{N}(v)\cup\{v\}}\frac{1}{\sqrt{\hat{d}_{u}\hat{d}_{v}}}\boldsymbol{h}_{u}^{l-1}\boldsymbol{W}^{l}))),$ (7)
@@ 4 Experimental Setup for Node Classification
Cora Homophily 2,708 5,278 2,708 7 Accuracy
@@
Squirrel Heterophily 5,201 216,933 2,089 5 Accuracy
Chameleon Heterophily 2,277 36,101 2,325 5 Accuracy
@@
GCN 85.08 ± 0.52 3.48$\uparrow$ 72.98 ± 0.84 1.38$\uparrow$ 81.32 ± 0.72 2.52$\uparrow$ 93.80 ± 0.29 4.15$\uparrow$ 96.51 ± 0.20 3.81$\uparrow$ 95.80 ± 0.28 2.88$\uparrow$ 97.43 ± 0.05 1.25$\uparrow$ 80.27 ± 0.71 2.80$\uparrow$
@@
GraphSAGE 84.18 ± 0.81 1.50$\uparrow$ 71.93 ± 0.85 0.00$\uparrow$ 79.41 ± 0.53 0.00$\uparrow$ 93.59 ± 0.22 2.39$\uparrow$ 96.41 ± 0.17 1.82$\uparrow$ 96.12 ± 0.24 2.21$\uparrow$ 97.21 ± 0.05 0.72$\uparrow$ 80.51 ± 0.48 5.74$\uparrow$
@@
GAT 84.64 ± 1.27 1.64$\uparrow$ 72.10 ± 1.10 0.00$\uparrow$ 79.70 ± 0.70 0.70$\uparrow$ 93.93 ± 0.16 3.15$\uparrow$ 96.67 ± 0.13 2.80$\uparrow$ 96.08 ± 0.10 2.47$\uparrow$ 97.30 ± 0.06 1.13$\uparrow$ 80.75 ± 0.74 3.84$\uparrow$
@@
Hyperparameter Configurations. We conduct minimal hyperparameter tuning on classic GNNs, consistent with the parameter search range of Polynormer [12]. We utilize the Adam optimizer [27] with a learning rate from $\{0.001,0.005,0.01\}$ and an epoch limit of 2500. And we tune the hidden dimension from $\{64,256,512\}$. As discussed in Section 3, we focus on whether to use normalization (BN or LN), residual connections, jumping knowledge, and dropout rates from $\{0.2,0.3,0.5,0.7\}$, the number of layers from $\{2,3,4,5,6,7,8,9,10\}$. Considering the large number of hyperparameters and datasets, we do not perform an exhaustive search. We report mean scores and standard deviations after 5 runs. GNN denotes our implementation of the GNN model.
@@ 5.1 Performance of Classic GNNs in Node Classification
GCN 44.50 ± 1.92 5.83$\uparrow$ 46.11 ± 3.16 4.80$\uparrow$ 53.57 ± 0.32 4.87$\uparrow$ 91.35 ± 0.37 17.68$\uparrow$ 97.77 ± 0.38 8.02$\uparrow$ 77.40 ± 1.07 1.31$\uparrow$
@@
GraphSAGE 39.93 ± 1.58 3.84$\uparrow$ 43.44 ± 3.19 5.67$\uparrow$ 54.72 ± 0.38 1.09$\uparrow$ 92.19 ± 0.58 6.45$\uparrow$ 96.95 ± 0.41 3.44$\uparrow$ 77.96 ± 0.72 1.52$\uparrow$
@@
GAT 38.72 ± 1.46 3.10$\uparrow$ 43.44 ± 3.00 4.23$\uparrow$ 54.99 ± 0.71 2.29$\uparrow$ 91.60 ± 0.21 2.85$\uparrow$ 97.76 ± 0.37 3.85$\uparrow$ 79.04 ± 1.27 2.25$\uparrow$
@@
GCN 77.29 ± 0.46 4.78$\uparrow$ 73.60 ± 0.18 1.86$\uparrow$ 82.33 ± 0.19 6.69$\uparrow$ 86.33 ± 0.17 10.88$\uparrow$
@@
GraphSAGE 82.21 ± 0.32 4.53$\uparrow$ 72.95 ± 0.31 1.46$\uparrow$ 83.89 ± 0.36 5.60$\uparrow$ 85.97 ± 0.21 10.34$\uparrow$
@@
GAT 85.01 ± 0.46 12.99$\uparrow$ 73.30 ± 0.16 1.35$\uparrow$ 80.99 ± 0.16 1.54$\uparrow$ 86.19 ± 0.23 13.96$\uparrow$
@@
While previously reported results show that most advanced GTs outperform classic GNN on homophilous graphs [12, 71], our implementation of classic GNNs can place within the top two for five datasets, with GCN always ranking in the top three. Specifically, on CS and WikiCS, classic GNNs experience about a 3% accuracy increase, achieving top-three performances. On WikiCS, the accuracy of GAT increases by 3.84%, moving it from seventh to first place, surpassing the leading GT, Polynormer. Similarly, on Computer and Photo, GAT outperforms Polynormer and SGFormer, establishing itself as the top model. On Cora, CiteSeer, PubMed, and Physics, tuning yields significant performance improvements for GCN, with accuracy increases ranging from 1.25% to 3.48%, positioning GCN as the highest-performing model despite its initial lower accuracy compared to advanced GTs.
@@
The three classic GNNs consistently secure top positions on five out of six heterophilous graphs. Specifically, on well-known page-page networks like Squirrel and Chameleon, our implementation enhances the accuracy of GCN by 4.80% and 5.83% respectively, elevating it to the first place among all models. By comparison, on larger heterophilous graphs such as Amazon-Ratings and Questions, GAT exhibits the highest performance, highlighting the superiority of its message-passing attention mechanism over GTs’ global attention. On Roman-Empire, a 17.68% increase is observed in the performance of GCN. Interestingly, we find that improvements primarily stem from residual connections, which are further analyzed in our ablation study (see Section 5.2).
@@
Our implementation of classic GNNs demonstrate consistently superior performance, achieving top rankings across all four large-scale datasets included in our study. Notably, GCN emerges as the leading model on ogbn-arxiv and pokec, surpassing all evaluated advanced GTs. Furthermore, on pokec, all three classic GNNs achieve over 10% performance increases by our implementation. For ogbn-proteins, an absolute improvement of 12.99% is observed in the performance of GAT, significantly surpassing SGFormer by 5.48%. Similarly, on ogbn-products, GraphSAGE demonstrates a performance increase, securing the best performance among all evaluated models. In summary, a basic GNN can achieve the best known results on large-scale graphs, suggesting that current GTs have not yet addressed GNN issues such as over-smoothing and long-range dependencies.
@@ 5.2 Influence of Hyperparameters on the Performance of GNNs
To examine the unique contributions of different hyperparameters in explaining the enhanced performance of classic GNNs, we conduct a series of ablation analysis by selectively removing elements such as normalization, dropout, residual connections, network depth, and jumping knowledge from GCN, GraphSAGE, and GAT. The effect of these ablations is assessed across homophilous (see Table 5), heterphilous (see Table 6), and large-scale graphs (see Table 7). Our findings, which we detail below, indicate that the ablation of single components affects model accuracy in distinct ways.
Table 5: Ablation study on homophilous graphs (%).
@@
GCN 85.08 ± 0.52 72.98 ± 0.84 81.32 ± 0.72 93.80 ± 0.29 96.51 ± 0.20 95.80 ± 0.28 97.43 ± 0.05 80.27 ± 0.71
(-) Normalization 85.08 ± 0.52 72.98 ± 0.84 81.32 ± 0.72 93.09 ± 0.13 96.38 ± 0.14 94.47 ± 0.25 97.18 ± 0.15 80.27 ± 0.71
(-) Dropout 82.70 ± 0.87 70.96 ± 1.00 79.50 ± 1.29 93.34 ± 0.11 96.51 ± 0.20 95.59 ± 0.16 97.32 ± 0.06 78.34 ± 0.42
(-) Residual Connections 83.40 ± 1.58 71.60 ± 1.26 80.86 ± 0.58 93.80 ± 0.29 94.68 ± 0.29 94.45 ± 0.12 96.89 ± 0.20 79.38 ± 0.57
(-) Jumping Knowledge 85.08 ± 0.52 72.98 ± 0.84 81.32 ± 0.72 92.97 ± 0.08 95.93 ± 0.32 93.73 ± 0.53 97.00 ± 0.10 79.13 ± 0.91
GraphSAGE 84.18 ± 0.81 71.93 ± 0.85 79.41 ± 0.53 93.59 ± 0.22 96.41 ± 0.17 96.12 ± 0.24 97.21 ± 0.05 80.51 ± 0.48
(-) Normalization 84.18 ± 0.81 71.80 ± 0.19 79.41 ± 0.53 93.07 ± 0.12 96.41 ± 0.17 94.53 ± 0.14 96.73 ± 0.05 80.41 ± 0.70
(-) Dropout 81.54 ± 1.55 70.36 ± 0.47 75.70 ± 1.69 92.71 ± 0.47 96.23 ± 0.10 96.12 ± 0.24 97.14 ± 0.04 80.51 ± 0.57
(-) Residual Connections 80.46 ± 1.19 71.80 ± 0.19 79.41 ± 0.53 93.23 ± 0.26 95.95 ± 0.13 95.87 ± 0.08 97.16 ± 0.04 80.00 ± 0.76
(-) Jumping Knowledge 84.18 ± 0.81 70.00 ± 1.26 79.41 ± 0.53 93.16 ± 0.18 95.82 ± 0.10 94.67 ± 0.26 96.91 ± 0.04 79.89 ± 0.79
GAT 84.64 ± 1.27 72.10 ± 1.10 79.70 ± 0.70 93.93 ± 0.16 96.67 ± 0.13 96.08 ± 0.10 97.30 ± 0.06 80.75 ± 0.74
(-) Normalization 84.64 ± 1.27 72.10 ± 1.10 79.70 ± 0.70 93.86 ± 0.10 96.21 ± 0.17 94.50 ± 0.33 96.99 ± 0.15 80.75 ± 0.74
(-) Dropout 83.18 ± 1.16 70.82 ± 0.76 79.04 ± 0.44 93.35 ± 0.33 96.67 ± 0.13 95.85 ± 0.04 97.30 ± 0.06 80.19 ± 0.16
(-) Residual Connections 83.34 ± 0.24 71.88 ± 0.40 79.70 ± 0.70 93.29 ± 0.31 95.29 ± 0.13 94.47 ± 0.25 96.51 ± 0.05 79.30 ± 0.92
(-) Jumping Knowledge 82.78 ± 0.96 70.36 ± 0.44 79.70 ± 0.70 93.44 ± 0.22 96.43 ± 0.20 94.41 ± 0.28 97.03 ± 0.04 79.56 ± 1.31
@@
GCN 44.50 ± 1.92 46.11 ± 3.16 53.57 ± 0.32 91.35 ± 0.37 97.77 ± 0.38 77.40 ± 1.07
(-) Normalization 43.92 ± 1.59 46.11 ± 3.16 53.57 ± 0.32 91.35 ± 0.37 97.77 ± 0.38 77.40 ± 1.07
(-) Dropout 41.96 ± 1.89 46.11 ± 3.16 50.99 ± 0.21 86.88 ± 0.34 94.33 ± 0.21 76.92 ± 1.55
(-) Residual Connections 44.50 ± 1.92 43.22 ± 3.91 50.31 ± 0.25 75.63 ± 0.48 89.16 ± 0.24 74.49 ± 0.39
(-) Jumping Knowledge 43.10 ± 1.47 43.81 ± 3.96 51.63 ± 0.27 90.73 ± 0.25 95.29 ± 0.62 77.30 ± 0.57
GraphSAGE 39.93 ± 1.58 43.44 ± 3.19 54.72 ± 0.38 92.19 ± 0.58 96.95 ± 0.41 77.96 ± 0.72
(-) Normalization 39.93 ± 1.58 43.44 ± 3.19 54.50 ± 0.38 92.19 ± 0.58 96.95 ± 0.41 76.07 ± 1.56
(-) Dropout 39.93 ± 1.58 40.98 ± 4.65 52.58 ± 0.17 86.50 ± 0.13 93.92 ± 0.73 77.96 ± 0.72
(-) Residual Connections 39.37 ± 2.30 41.24 ± 3.85 53.94 ± 0.40 88.65 ± 0.96 94.41 ± 0.43 76.70 ± 0.87
(-) Jumping Knowledge 38.70 ± 1.95 43.44 ± 3.19 54.27 ± 0.69 90.12 ± 0.54 93.80 ± 0.30 76.80 ± 0.72
GAT 38.72 ± 1.46 43.44 ± 3.00 54.99 ± 0.71 91.60 ± 0.21 97.76 ± 0.37 79.04 ± 1.27
(-) Normalization 38.72 ± 1.46 42.24 ± 3.26 54.99 ± 0.71 90.96 ± 0.20 97.76 ± 0.37 79.02 ± 1.13
(-) Dropout 38.72 ± 1.46 41.08 ± 2.96 53.84 ± 0.87 87.06 ± 0.21 94.10 ± 0.09 78.51 ± 0.36
(-) Residual Connections 38.72 ± 1.46 41.74 ± 2.08 51.24 ± 0.28 79.11 ± 0.88 90.83 ± 0.45 77.07 ± 0.84
(-) Jumping Knowledge 38.72 ± 1.46 41.12 ± 3.94 54.38 ± 0.29 91.60 ± 0.21 93.92 ± 0.73 79.04 ± 1.27

### v2

@@ Abstract
Graph Transformers (GTs) have recently emerged as popular alternatives to traditional message-passing Graph Neural Networks (GNNs), due to their theoretically superior expressiveness and impressive performance reported on standard node classification benchmarks, often significantly outperforming GNNs. In this paper, we conduct a thorough empirical analysis to reevaluate the performance of three classic GNN models (GCN, GAT, and GraphSAGE) against GTs. Our findings suggest that the previously reported superiority of GTs may have been overstated due to suboptimal hyperparameter configurations in GNNs. Remarkably, with slight hyperparameter tuning, these classic GNN models achieve state-of-the-art performance, matching or even exceeding that of recent GTs across 17 out of the 18 diverse datasets examined. Additionally, we conduct detailed ablation studies to investigate the influence of various GNN configurations—such as normalization, dropout, residual connections, and network depth—on node classification performance. Our study aims to promote a higher standard of empirical rigor in the field of graph machine learning, encouraging more accurate comparisons and evaluations of model capabilities. Our implementation is available at https://github.com/LUOyk1999/tunedGNN.
@@ 1 Introduction
Node classification is a fundamental task in graph machine learning [92, 79, 54, 78, 71, 77], with high-impact applications across many fields such as social network analysis, bioinformatics, and recommendation systems. Graph Neural Networks (GNNs)  [20, 28, 69, 80, 52, 33, 8, 84, 18, 9, 55, 4, 81, 60, 10, 56, 57, 70, 85, 37] have emerged as a powerful class of models for tackling the node classification task. GNNs operate by iteratively aggregating information from a node’s neighbors, a process known as message passing [19], leveraging both the graph structure and node features to learn useful node representations for classification. While GNNs have achieved notable success, studies have identified several limitations, including over-smoothing [35], over-squashing [1], lack of sensitivity to heterophily [90], and challenges in capturing long-range dependencies [11].
@@
This prompts us to reconsider: Could the potential of message-passing GNNs for node classification have been previously underestimated? While prior research has addressed this issue to some extent [24, 14, 73, 47, 58], these studies have limitations in terms of scope and comprehensiveness, including a restricted number and diversity of datasets, as well as an incomplete examination of hyperparameters. In this study, we comprehensively reassess the performance of GNNs for node classification, utilizing three classic GNN models—GCN [28], GAT [68], and GraphSAGE [20]—across 18 real-world benchmark datasets that include homophilous, heterophilous, and large-scale graphs. We examine the influence of key hyperparameters on GNN training, including normalization [2, 26], dropout [67], residual connections [21], and network depth. We summarize the key findings in our empirical study as follows:
@@
  • Our ablation studies have yielded valuable insights into GNN hyperparameters for node classification. We demonstrate that (1) normalization is essential for large-scale graphs; (2) dropout consistently proves beneficial; (3) residual connections can significantly enhance performance, especially on heterophilous graphs; and (4) GNNs on heterophilous graphs tend to perform better with deeper layers.
@@ 3 Key Hyperparameters for Training GNNs
In this section, we present an overview of the key hyperparameters for training GNNs, including normalization, dropout, residual connections, and network depth. These hyperparameters are widely utilized across different types of neural networks to improve model performance.
@@
$\boldsymbol{h}_{v}^{l}=\text{Dropout}(\sigma(\text{Norm}(\boldsymbol{h}_{v}^{l-1}\boldsymbol{W_{r}}^{l}+\sum_{u\in\mathcal{N}(v)\cup\{v\}}\frac{1}{\sqrt{\hat{d}_{u}\hat{d}_{v}}}\boldsymbol{h}_{u}^{l-1}\boldsymbol{W}^{l}))),$ (7)
@@ 4 Experimental Setup for Node Classification
Cora Homophily 2,708 5,278 1,433 7 Accuracy
@@
Squirrel Heterophily 2,223 46,998 2,089 5 Accuracy
Chameleon Heterophily 890 8,854 2,325 5 Accuracy
@@
GCN 85.10 ± 0.67 3.50$\uparrow$ 73.14 ± 0.67 1.54$\uparrow$ 81.12 ± 0.52 2.32$\uparrow$ 93.99 ± 0.12 4.34$\uparrow$ 96.10 ± 0.46 3.40$\uparrow$ 96.17 ± 0.06 3.25$\uparrow$ 97.46 ± 0.10 1.28$\uparrow$ 80.30 ± 0.62 2.83$\uparrow$
@@
GraphSAGE 83.88 ± 0.65 1.20$\uparrow$ 72.26 ± 0.55 0.33$\uparrow$ 79.72 ± 0.50 0.31$\uparrow$ 93.25 ± 0.14 2.05$\uparrow$ 96.78 ± 0.23 2.19$\uparrow$ 96.38 ± 0.11 2.47$\uparrow$ 97.19 ± 0.05 0.70$\uparrow$ 80.69 ± 0.31 5.92$\uparrow$
@@
GAT 84.46 ± 0.55 1.46$\uparrow$ 72.22 ± 0.84 0.12$\uparrow$ 80.28 ± 0.64 1.28$\uparrow$ 94.09 ± 0.37 3.31$\uparrow$ 96.60 ± 0.33 2.73$\uparrow$ 96.21 ± 0.14 2.60$\uparrow$ 97.25 ± 0.06 1.08$\uparrow$ 81.07 ± 0.54 4.16$\uparrow$
@@
Hyperparameter Configurations. We conduct hyperparameter tuning on classic GNNs, consistent with the hyperparameter search space of Polynormer [12]. Specifically, we utilize the Adam optimizer [27] with a learning rate from $\{0.001,0.005,0.01\}$ and an epoch limit of 2500. And we tune the hidden dimension from $\{64,256,512\}$. As discussed in Section 3, we focus on whether to use normalization (BN or LN), residual connections, and dropout rates from $\{0.2,0.3,0.5,0.7\}$, the number of layers from $\{1,2,3,4,5,6,7,8,9,10\}$. Additionally, we retrain all baseline GTs using the same hyperparameter search space and training environments as the classic GNNs. For hyperparameters specific to each GT, which are not present in the classic GNNs, we tune them according to the search space specified in the original GT paper. We report mean scores and standard deviations after 5 independent runs with different initializations. Model denotes our implementation.
@@ 5.1 Performance of Classic GNNs in Node Classification
GCN 45.01 ± 1.63 6.34$\uparrow$ 46.29 ± 3.40 4.98$\uparrow$ 53.80 ± 0.60 5.10$\uparrow$ 91.27 ± 0.20 17.58$\uparrow$ 97.86 ± 0.24 8.11$\uparrow$ 79.02 ± 0.60 2.93$\uparrow$
@@
GraphSAGE 40.78 ± 1.47 4.69$\uparrow$ 44.81 ± 4.74 7.04$\uparrow$ 55.40 ± 0.21 1.77$\uparrow$ 91.06 ± 0.27 5.32$\uparrow$ 97.77 ± 0.62 4.26$\uparrow$ 77.21 ± 1.28 0.77$\uparrow$
@@
GAT 41.73 ± 2.07 6.11$\uparrow$ 44.13 ± 4.17 4.92$\uparrow$ 55.54 ± 0.51 2.84$\uparrow$ 90.63 ± 0.14 1.88$\uparrow$ 97.73 ± 0.73 3.82$\uparrow$ 77.95 ± 0.51 1.16$\uparrow$
@@
GCN 77.29 ± 0.46 4.78$\uparrow$ 73.53 ± 0.12 1.79$\uparrow$ 82.33 ± 0.19 6.69$\uparrow$ 86.33 ± 0.17 10.88$\uparrow$
@@
GraphSAGE 82.21 ± 0.32 4.53$\uparrow$ 73.00 ± 0.28 1.51$\uparrow$ 83.89 ± 0.36 5.60$\uparrow$ 85.97 ± 0.21 10.34$\uparrow$
@@
GAT 85.01 ± 0.46 12.99$\uparrow$ 73.30 ± 0.18 1.35$\uparrow$ 80.99 ± 0.16 1.54$\uparrow$ 86.19 ± 0.23 13.96$\uparrow$
@@
While previously reported results show that most advanced GTs outperform classic GNN on homophilous graphs [12, 76], our implementation of classic GNNs can place within the top two for four datasets, with GCN and GAT demonstrating near-consistent top performances. Specifically, on CS and WikiCS, classic GNNs experience about a 3% accuracy increase, achieving top-three performances. On WikiCS, the accuracy of GAT increases by 4.16%, moving it from seventh to first place, surpassing the leading GT, Polynormer. Similarly, on Photo and CS, GraphSAGE outperforms Polynormer and SGFormer, establishing itself as the top model. On Cora, CiteSeer, PubMed, and Physics, tuning yields significant performance improvements for GCN, with accuracy increases ranging from 1.54% to 3.50%, positioning GCN as the highest-performing model despite its initial lower accuracy compared to advanced GTs.
@@
The three classic GNNs secure top positions on five out of six heterophilous graphs. Specifically, on well-known page-page networks like Chameleon and Squirrel, our implementation enhances the accuracy of GCN by 4.98% and 6.34% respectively, elevating it to the first place among all models. Similarly, on larger heterophilous graphs such as Minesweeper and Questions, GCN also exhibits the highest performance, highlighting the superiority of its local message-passing mechanism over GTs’ global attention. On Roman-Empire, a 17.58% increase is observed in the performance of GCN. Interestingly, we find that improvements primarily stem from residual connections, which are further analyzed in our ablation study (see Section 5.2).
@@
Our implementation of classic GNNs demonstrate superior performance consistently, achieving top rankings across all four large-scale datasets included in our study. Notably, GCN emerges as the leading model on ogbn-arxiv and pokec, surpassing all evaluated advanced GTs. Furthermore, on pokec, all three classic GNNs achieve over 10% performance increases by our implementation. For ogbn-proteins, an absolute improvement of 12.99% is observed in the performance of GAT, significantly surpassing SGFormer by 5.09%. Similarly, on ogbn-products, GraphSAGE demonstrates a significant performance increase, securing the best performance among all evaluated models. In summary, a basic GNN can achieve the best known results on large-scale graphs, suggesting that current GTs have not yet addressed GNN issues such as over-smoothing and long-range dependencies.
@@ 5.2 Influence of Hyperparameters on the Performance of GNNs
To examine the unique contributions of different hyperparameters in explaining the enhanced performance of classic GNNs, we conduct a series of ablation analysis by selectively removing elements such as normalization, dropout, residual connections, and network depth from GCN, GraphSAGE, and GAT. The effect of these ablations is assessed across homophilous (see Table 5), heterphilous (see Table 6), and large-scale graphs (see Table 7). Our findings, which we detail below, indicate that the ablation of single components affects model accuracy in distinct ways.
Table 5: Ablation study on homophilous graphs (%). - indicates that the corresponding hyperparameter is not used in GNN, as it empirically leads to inferior performance.
@@
GCN 85.10 ± 0.67 73.14 ± 0.67 81.12 ± 0.52 93.99 ± 0.12 96.10 ± 0.46 96.17 ± 0.06 97.46 ± 0.10 80.30 ± 0.62
(-) Normalization - - - 92.60 ± 0.14 95.48 ± 0.36 95.30 ± 0.05 97.16 ± 0.11 79.67 ± 0.52
(-) Dropout 83.46 ± 1.16 71.40 ± 0.35 80.14 ± 0.55 93.78 ± 0.26 95.31 ± 0.10 95.95 ± 0.14 97.30 ± 0.06 79.84 ± 0.86
(-) Residual Connections - - - - 94.43 ± 0.16 94.71 ± 0.12 96.56 ± 0.18 -
GraphSAGE 83.88 ± 0.65 72.26 ± 0.55 79.72 ± 0.50 93.25 ± 0.14 96.78 ± 0.23 96.38 ± 0.11 97.19 ± 0.05 80.69 ± 0.31
(-) Normalization - - - 92.77 ± 0.63 95.51 ± 0.46 95.42 ± 0.13 96.97 ± 0.07 80.08 ± 0.85
(-) Dropout 82.78 ± 0.54 71.02 ± 1.34 77.02 ± 0.83 92.02 ± 0.35 96.03 ± 0.27 96.11 ± 0.17 97.07 ± 0.09 79.89 ± 0.39
(-) Residual Connections - - - - 96.47 ± 0.11 95.73 ± 0.13 97.09 ± 0.04 -
GAT 84.46 ± 0.55 72.22 ± 0.84 80.28 ± 0.64 94.09 ± 0.37 96.60 ± 0.33 96.21 ± 0.14 97.25 ± 0.06 81.07 ± 0.54
(-) Normalization - - - 93.22 ± 1.27 96.09 ± 0.20 95.13 ± 0.16 97.08 ± 0.04 80.34 ± 0.41
(-) Dropout 83.30 ± 1.34 71.22 ± 1.02 78.36 ± 1.22 93.14 ± 0.29 96.36 ± 0.26 96.05 ± 0.09 97.01 ± 0.05 79.46 ± 0.32
(-) Residual Connections 82.62 ± 1.08 71.60 ± 0.89 - - 95.05 ± 0.14 94.47 ± 0.09 96.44 ± 0.03 80.32 ± 0.90
@@
GCN 45.01 ± 1.63 46.29 ± 3.40 53.80 ± 0.60 91.27 ± 0.20 97.86 ± 0.24 79.02 ± 0.60
(-) Normalization 44.13 ± 2.03 - 53.68 ± 0.82 90.53 ± 0.33 96.94 ± 1.96 -
(-) Dropout 42.89 ± 1.28 45.28 ± 4.78 51.37 ± 0.34 85.10 ± 0.61 94.28 ± 2.29 76.58 ± 0.40
(-) Residual Connections 43.14 ± 1.82 - 51.14 ± 0.34 74.84 ± 0.62 86.45 ± 0.89 75.87 ± 4.47
GraphSAGE 40.78 ± 1.47 44.81 ± 4.74 55.40 ± 0.21 91.06 ± 0.27 97.77 ± 0.62 77.21 ± 1.28
(-) Normalization 40.27 ± 2.27 44.02 ± 3.53 54.41 ± 0.30 90.58 ± 0.24 97.64 ± 0.41 76.17 ± 0.41
(-) Dropout 38.83 ± 1.94 43.11 ± 3.36 51.12 ± 0.66 84.49 ± 0.35 93.83 ± 0.38 76.36 ± 1.50
(-) Residual Connections 40.06 ± 2.31 41.85 ± 3.86 53.52 ± 0.19 - 96.64 ± 0.85 -
GAT 41.73 ± 2.07 44.13 ± 4.17 55.54 ± 0.51 90.63 ± 0.14 97.73 ± 0.73 77.95 ± 0.51
(-) Normalization 41.08 ± 1.63 43.25 ± 3.84 54.85 ± 0.39 89.69 ± 0.39 97.42 ± 0.85 76.32 ± 0.24
(-) Dropout 39.81 ± 3.15 41.19 ± 2.36 51.48 ± 0.28 82.47 ± 0.70 92.26 ± 4.63 76.19 ± 0.88
(-) Residual Connections 38.46 ± 1.96 42.57 ± 3.66 51.08 ± 0.49 85.15 ± 0.82 92.83 ± 1.61 75.17 ± 0.71